OpenStudy (anonymous):

HELP PLEASE!!! What is the solution set of the equation 2 cos^2 data- cos data =0 over the interval 0 degrees < or equal to data < 360 degrees?

Miracrown (miracrown):

Not sure why the variable is called "data" but its the same as: \[2\cos^2x \space - cosx = 0 \space \space , 0 \le x <360\]

Miracrown (miracrown):

right? This is the same equation? So we could solve this with factoring, do you see how to proceed in that direction?

OpenStudy (anonymous):

nope, not at all

OpenStudy (anonymous):

btw the answer is {60, 90, 270, 300} degrees

OpenStudy (anonymous):

and again, please help me still, cause i don't don't know how the heck to do, please

Miracrown (miracrown):

Is there a GCF?

Miracrown (miracrown):

cosx is common right?

OpenStudy (anonymous):

uh cosx?

OpenStudy (anonymous):

oh well I guess we both answered it then lol

OpenStudy (anonymous):

ok so please continue....

Miracrown (miracrown):

\[cosx(2cosx-1)=0\]

Miracrown (miracrown):

then just set each factor equal to 0 and see what solutions result

OpenStudy (anonymous):

I still don't get how the answer is 60, 90, 270, and 300

Miracrown (miracrown):

\[cosx = 0\] \[2cosx-1=0\]

OpenStudy (anonymous):

I don't know what to do from there

Miracrown (miracrown):

Well, are you familiar with the Unit Circle?

OpenStudy (anonymous):

yes

Miracrown (miracrown):



Miracrown (miracrown):

so we solve those 2 equations for cosx first

Miracrown (miracrown):

we want to find the angles where cosx=0 and cosx= ..?

OpenStudy (anonymous):

0?

Miracrown (miracrown):

first is cosx=0 2nd is different 2cosx-1 = 0 2cosx = 1

Miracrown (miracrown):

so its cosx = 1/2 right?

Miracrown (miracrown):

So we want to find on the Unit circle , the angles that have a cosine of 0, or 1/2 On the Unit circle is it the x, or the y, that corresponds to cosine?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

wait wait wait wait wait

OpenStudy (anonymous):

it's 1/2 if it's pie over 3 if it's pie over 6, sin is 1/2

Miracrown (miracrown):

The terminal point on the Unit Circle is (cos(theta), sin(theta)) So its the x so I want to see what point has an x=1/2 or an x=0 So I can draw 2 vertical lines x=0 and x=1/2 so where they intersect the unit circle will be our solutions. But we still have to have a good idea what angles we are looking at

Miracrown (miracrown):

yes so that is pi/3 or 60 deg the first one

OpenStudy (anonymous):

wait how u know it's 1/2 again, that's the part i don't get?

OpenStudy (anonymous):

cause I thought it could either be 1/2 or radical 3/2 right???

Miracrown (miracrown):

because our solution said cosx = 1/2

Miracrown (miracrown):

on the unit circle that means the x-coordinate of the terminal point is 1/2

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh LOL

OpenStudy (anonymous):

I'm so dumb I don't know how to add LOL!!!

Miracrown (miracrown):

Haha

Miracrown (miracrown):

http://prntscr.com/6bnn3e

OpenStudy (anonymous):

ok please continue

Miracrown (miracrown):

on the unit circle that means the x-coordinate of the terminal point is 1/2 so I drew a vertical line x = 1/2 and looked to see where it intersected the Unit Circle and of course I know it will be some multiple of 30 deg I just have to figure out which qudrant and then if in Quad I, is it 30 or 60? and so on

OpenStudy (anonymous):

ok so what about 90, 270, and 300???